OpenStudy (anonymous):

If f(x) = x^2 + 3, find f(5). A) 13 B) 22 C) 28 D) 4

jimthompson5910 (jim_thompson5910):

f(x) = x^2 + 3 f(5) = 5^2 + 3 f(5) = 25 + 3 f(5) = 28

OpenStudy (anonymous):

f(5) = (5)^2 + 3 = 28

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Thanks:)